• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Epstein, templon

    Log entry time 09:33:12 on December 18,1999

    Entry number 31391

    keyword=e spect change, new espace hdr

    Made a new header file for ESPACE for kin. # 17, fixed the field setting for the electron spectrometer. New header file is: hdr_he3_17. this header
    file solved the problems we had with Emiss = -440 MeV etc

    we appear to now have about 90k events in the 3He(e,e'p)2H peak
    so we stop this kin. point and will begin the transition to kinematics
    27. the first step is to change the electron arm from 95.94 deg to
    87.40 deg, assoc. momentum setting 583.19 MeV/c.